How many ways can a row of four girls followed by four boys be arranged if the
girls must sit together and the boys must sit together?
(A) 16 (B) 256 (C) 576 (D) 10,080 (E) 40,320

Answer by amoresroy(361)

How many ways can a row of four girls followed by four boys be arranged if the
girls must sit together and the boys must sit together?
(A) 16 (B) 256 (C) 576 (D) 10,080 (E) 40,320
Answer: (C) 576
Solution:
(4*3*2*1)*(4*3*2*1)= 24*24 = 576